OpenStudy (anonymous):

Andy has grades of 84, 65, and 76 on three math tests. What grade must he obtain on the next test to have an average of exactly 80 for the four tests?

OpenStudy (anonymous):

(84+65+76+x)/4=80 (225+x)/4=80 225+x=320 x=95

OpenStudy (anonymous):

So he needs to score a 95% on his next test.

OpenStudy (anonymous):

Hope this helped!

OpenStudy (anonymous):

Thanks so much|

OpenStudy (anonymous):

Anytime.

OpenStudy (anonymous):

can you show the work please

OpenStudy (anonymous):

I did

OpenStudy (anonymous):

Look at my first reply.

OpenStudy (anonymous):

lol can you do it more neater

OpenStudy (anonymous):

Absolutly

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

The average of four test scores is the sum of the 4 divided by 4, so: (a + b + c + d) / 4 We have 3 of the 4 scores and want the average to be 80, so: 80 = (84 + 65 + 76 + d) / 4 Solve for d: 320 = 84 + 65 + 76 + d 320 = 225 + d 95 = d The fourth test score must be a 95 to get an average of 80.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

i still dont unnderstand :'

OpenStudy (anonymous):

sadly

OpenStudy (anonymous):

hmmmmm, I will try to explain it better.

OpenStudy (anonymous):

thankss

OpenStudy (anonymous):

Anytime

OpenStudy (anonymous):

84 is 80 + 4 65 is 80 - 15 76 is 80 - 4 Overall, he is 15 points below 80 over three tests. He has to make up 15 points. If he makes a 95, his average will be 80.

OpenStudy (anonymous):

Better?

OpenStudy (anonymous):

omg thanks this helped alot

OpenStudy (anonymous):

haha, no problem :)

OpenStudy (anonymous):

wait but why is it only 3 numbers

OpenStudy (anonymous):

Because the fourth one is the one he needs to get

OpenStudy (anonymous):

He has taken 3 tests and needs one more to get the given average.

OpenStudy (anonymous):

oh okayy

OpenStudy (anonymous):

Well, I have to go help others now, but judging by this problem your in 7th grade right?